OpenStudy (goldrush18):

the roots of the equation 2x^2-4x+7=0 are alpha and beta. find the value of the roots of the equation 2x^2-4x+7=0 are alpha and beta. find the value of @Mathematics

OpenStudy (goldrush18):

\[1/\alpha+2 + 1/\beta+2\]

OpenStudy (goldrush18):

and also \[\alpha^2\beta+\beta^2\alpha\]